Using values from Appendix C of your textbook, calculate the value of Keq at 298 K for each of the following reactions:Using values from Appendix C of your textbook, calculate the value of Keg at 298 K for each of the following reactions: (a) F
Answer #1

Ans. Güven Peaction- @ Fe3O4 (s) + 4C0C9? = 46° = -2.303 RT logk 3 Fe (s) + 460968) AG° - Standareel free energy R = Cros. Co= [1547.6 kJ molt ] - [1019.2 ks mort+(-548-8 k smi] = -1577.6 - [ - 1563] = (1577.6 + 1563) K J molt = -14.6 kJ mort AG° = -46° = [ax 6334 kImalt) + 2 * (-228-57 Kumar)] - [-740.98 kJ molt + 6x-95-27 ks mi] = [668 - 457.14] - [ -740 – 571.62] = (-1461 = [2 x 444°C HgHq) + 2x 446°C4206)] - [ 4x Asci°C NH3] + Af6° 003] = (2x159.4 kJmort + 2x(-237.13 kJmort] - [4X (-16.6 ks